OpenStudy (anonymous):

simplify: 1/9 - 1/x divided by 1/81 - 1/x^2 answers a)9/x+9 b)x+9/9 c)x+9/9x d)9x/x+9

OpenStudy (anonymous):

D.

OpenStudy (anonymous):

can you show me how you worked that? please.

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

(1/81 - 1/x^2) = (1/9)^2 - (1/x)^2 = (1/9 - 1/x)*(1/9 + 1/x)

jhonyy9 (jhonyy9):

1/9 -1/x =(x-9)/9x 1/81 - 1/x^2 =(x^2 -81)/81x^2 =(x-9)(x+9)/81x^2 (x-9) ----- 9x (x-9) 81x^2 ------------ = ------- * ----------- = 9x/(x+9) so that is D. (x-9)(x+9) 9x (x-9)(x+9) ----------- 81x^2

OpenStudy (anonymous):

(1/9 - 1/x) / (1/81 - 1/x^2)

OpenStudy (anonymous):

yeah like what jhony said lol

OpenStudy (anonymous):

thanks you guys!!

jhonyy9 (jhonyy9):

good luck bye

OpenStudy (anonymous):

:)